Sol Summary – January 28th

Sol: 7
Summary Title: A day of rest
Author’s name: Andrew Wheeler
Mission Status: Nominal
Sol Activity Summary: Today was scheduled to be a day of rest and personal time. However, it didn’t quite turn out that way. One crew member awoke to a severe migraine and another to a stomach upset and their wellbeing was our main concern. Nevertheless, washing and cleaning was able to be done and a very short EVA to return the solar logger it’s recording point was carried out in the early morning. The Mars jigsaw puzzle was completed and signed and an issue with the lower level gas heater was remedied.
Look Ahead Plan: We will return to robot rover testing on the ground around the hab in the morning and an EVA to the stream draining from Tharsis Montes in search of concretions in the afternoon. If the weather holds, the solar observatory will also be operated.
Anomalies in work: Lower level gas heater anomaly – see operations report.
Weather: Clear crisp morning at -5 Celsius rising to low teens during the day. High cirrus cloud later in the day. No precipitation. Humidity continuously dropping from near 60% to below 40%. The wind was faint to gentle during the day.
Crew Physical Status: Severe migraine headache as well as stomach upset – see HSO report.
EVA: EVA 12 with Andrew and Clare to return the solar logger to it’s recording position and an external inspection of the hab tunnels.
Reports to be filed: Sol Summary, Commander’s Report, Greenhab Report, Operations Report (including an HSO report), EVA report, EVA Request(s). There is no Astronomer report for today.
Support Requested: Nothing to request.

Copyright © The Mars Society. All rights reserved. | Main Site